OpenStudy (imperfection):

Mr.Jones is 30 years older than his son. Six years ago he was seven times as old as his son was then. Find their present ages.

OpenStudy (anonymous):

I think Mr.Jones is 33, and his son is 4? I'm not entirely sure...?

OpenStudy (imperfection):

i need steps...

OpenStudy (imperfection):

i got 11 and 41.

OpenStudy (anonymous):

First take 30-6.. That gives you 24! Now divide 24 by 7 and you get 4, correct? Yes, so his son is 4, and add 3 years to Mr.Jones, and then add 24 plus the 7 and age of 4..?

OpenStudy (imperfection):

x+30-6=7(x-6) x+24=7x-42 6x=66 x=11

OpenStudy (anonymous):

I guess?

OpenStudy (imperfection):

ok thank you. i got it

OpenStudy (anonymous):

Your welcome:DDD